• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:15:30 on November 9, 1998

    Entry number 5025

    keyword=Beam Time Accounting
    Hall A

    11/9/1998 0800-1600
    Experiment #: E94-010
    Shift Leader: B. Hersman
    Shift Workers: E. Chudakov, C. Jutier, S. Incerti, R. Suleiman

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    1 1 0.9 0.5 1 1 1 .9 7.3
    Beam Available
    but Not in Use
    0 0 0 0 0 0 0 0 0
    Beam Not Available
    or Unacceptable
    0 0 0.1 0.5 0 0 0 .1 0.7
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 1 1 1 1 1 1 8
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: NMR measurement gives 29.3% after completing transverse measurement. We will pump in longitudinal configuration while the Moller and e-p measurements are in progress, and measure again before beginning longitudinal running. - Lasers off at 9:31, to allow temperature stabilization for 15 minutes. We plan to take a "cold" NMR measurement, followed by a "warm" NMR measurement (lasers ON). - Moeller : Conditions: 35 degrees on the injector, 4.255GeV. The beam polarization measured is 72.9+/-0.2(stat)+/-3.0(prel_syst)%. - NMR measurement with lasers off (and cooled down) gives 28.36% Moller also complete. Preparing for e-p with 2.5 uA and raster requested OFF - NMR at 10:54 AM : 29.88 % (H2O) and 33.89% (EPR). Parallel configuration - no beam since 10:55. Pblms to set up fast raster. - ep data taking by Yves from 9:51 to 10:41 AM - raster checked (r=1.75 mm). Runs at 1, 5 , 10 and 15 uA

    Data Entry by: B Hersman